The sum of two numbers is 90.  If three times the smaller number is subtracted from the larger​ number, the result is 14.  Find the two numbers.
Answer by ikleyn(52835)

.
Let x be the smaller number.

Then the greater number is (90-x).


The condition then says


    (90-x) - 3x = 14.


Simplify and solve for x.


    90 - x - 3x = 14

    90 - 14 = x + 3x

    76 = 4x  =================>  x = 76/4 = 19.


ANSWER.  The smaller number is 19;  the larger number is 90-19 = 71.


CHECK.   71 - 3*19 = 71 - 57 = 14.     ! Correct !

Solved.